# DriveMLM: Aligning Multi-Modal Large Language Models with Behavioral Planning States for Autonomous Driving

### v2

@@ 4.5 Ablation Study
Sensor Modality
Table 5 presents the results of different impacts of input sensor modality to the DriveMLM. Multi-View (MV) images bring a substantial performance improvement in both path and speed F1-score, along with 18.19% increase in accuracy. Compared to concatenating temporal tokens directly, temporal QFormer results in a larger improvement of 7.4%, while ensuring multi-modal decision capability, which leads to 0.05 improvement in the average F1-score on speed decision. Point clouds do not show the ability to enhance performance.
[…]
@@
Table 5: Ablation results of sensor modality and temporal information. MV denotes multi-view images, CT denotes concatenating temporal tokens, TQ denotes temporal QFormer, and PC denotes point clouds. MV + TQ shows the best decision performance, and CT brings a small improvement in accuracy but leads to greater computational consumption. PC has little impact on DriveMLM. This might caused by the large representation gap between the sparse pyramid transformer and the MLLM Decoder.
Temporal Module Design
We propose the temporal QFormer module to process the temporal multi-view images. A simple and naive design is directly concatenating query tokens temporal to generate $N_{tq}=T\times N_{I}\times N_{Q}$ tokens acting as LLM input. But $N_{tq}$ increases with $T$, contributing to large computational costs. Instead, we propose the temporal QFormer module to process temporal images for each view separately, generating $N_{I}\times N_{Q}$ tokens for LLM input. The comparison of the temporal module is shown in table 5, indicating the better performance of our temporal module design with fewer image tokens. We set $T=2$ by default in our experiments.
@@
4.6 Case Study & Visualization
Human Interaction
Figure 4 provides an example of how vehicle control can be achieved through human instructions. The control process involves analyzing the road conditions, making decision choices, and providing explanatory statements. When given the identical instruction to “overtake”, DriveMLM exhibits varying responses based on the analysis of the current traffic conditions. In the scenario where the right lane is occupied and the left lane is available, the system opts to overtake from the left. However, in situations where the given instruction may pose a danger, such as when all lanes are occupied, DriveMLM chooses to refrain from executing the overtaking maneuver and responds appropriately. DriveMLM, in this context, serves as an interface for human-vehicle interaction, which evaluates the reasonableness of the instruction based on traffic dynamics and ensures its compliance with predefined rules before ultimately selecting a course of action.
Performance in Real Scenarios
We apply DriveMLM on the nuScenes dataset [8] to test the zero-shot performance of the developed driving system. We annotate 6,019 frames on the validation set, and the zero-shot performance of decision accuracy is 0.395. Figure 5 presents the result on two real driving scenes, indicating the generability of DriveMLM.
@@ 5 Conclusion
[…]
Supplementary Material
Appendix A Prompt Details
As illustrated in Table A, we provide the complete system message which includes detailed definitions of path decision states and speed decision states. Specifically, our path decision states include 5 states, which are {FOLLOW_LANE, LEFT_LANE_CHANGE, RIGHT_LANE_CHANGE, LEFT_LANE_BORROW, RIGHT_LANE_BORROW}, and our speed decision states include 4 states: {KEEP, ACCELERATE, DECELERATE, STOP}.
In Table B, we detail the prompts utilized for describing the surrounding environments. In Table C, prompts are employed to derive driving decisions based on navigation commands. The prompts listed in Table D are used to elicit explanations from the model regarding its decisions. Finally, in Table E, we present the human instructions that were used to guide the model’s decision-making process.
System Message: You are a driving assistant to drive the car. You need to follow the navigation command and traffic rules. The traffic rule is 1. Traffic light indications: a. Green: Vehicles may proceed. b. Yellow: Vehicles already past the stop line can continue. c. Red: Vehicles must stop. 2. Vehicle regulations: a. Vehicles must not exceed speed limits indicated by signs or road markings. b. Vehicles must stop when they meet the stop line. 3. Drivers should note specific traffic signs/markings: - Double solid lines: Overtaking is prohibited. Adhere strictly and don’t cross to overtake. - Single solid line: Overtaking is restricted. Overtaking is allowed to provide a safe distance and clear visibility, ensuring safety. 4. If special vehicles like police or ambulances are behind, yield and allow them to pass first. 5. Collision with other moving or static objects is not allowed. Path decision definitions: ‘LEFT_LANE_CHANGE’ refers to a driver’s decision to switch from the current to the adjacent left lane. ‘RIGHT_LANE_CHANGE’ refers to a driver’s decision to switch from the current lane to the adjacent right lane. ‘LEFT_LANE_BORROW’ is when a driver temporarily uses the adjacent left lane, commonly for overtaking or avoiding obstacles. ‘RIGHT_LANE_BORROW’ is when a driver temporarily uses the adjacent right lane, commonly for overtaking or avoiding obstacles. ‘FOLLOW_LANE’ means the driver decides to continue in their current lane. Speed decision definitions: ‘ACCELERATE’ refers to a driver increasing their speed. ‘DECELERATE’ means the driver reduces their speed. ‘KEEP’ refers to a driver keeping a steady speed. ‘STOP’ means the driver completely halts the vehicle. Based on the definitions of path decision, and while adhering to traffic rules, please choose a path and speed decision from the predefined options below, considering the current scenario. Path decisions include [LEFT_LANE_BORROW, RIGHT_LANE_BORROW, LEFT_LANE_CHANGE, RIGHT_LANE_CHANGE, FOLLOW_LANE]. Speed decisions include [ACCELERATE, DECELERATE, KEEP, STOP]. Given the navigation command and driving scene obtained from the camera or LiDAR, You should choose a path decision and a speed decision from the predefined options and give an explanation of your decision.
Table A: System message. The system message includes the description of the driving task, the traffic rules, and the definition of decision states.
1. This is a driving scenario. Please describe the environment. The images are provided by the front, left, right, and back cameras of a vehicle. The point cloud is generated by a LiDAR mounted on the top of the vehicle. 2. Could you provide me with a description of the current surroundings? Visual information from the vehicle’s front, left, right, and back cameras provides the images, while a LiDAR sensor mounted on the top of the vehicle generates a point cloud representation of the environment. 3. Would you kindly provide me with a description of the current surroundings? The vehicle’s front, left, right, and back cameras provide the visual images, while a LiDAR system, mounted on the top of the vehicle, generates a point cloud representation of the environment. 4. Could you please describe the current surroundings to me? 5. Could you kindly provide me with a description of the surrounding environment, please?
Table B: Prompts used to describe the surrounding environment.
1. The navigation command is turn left. Please choose a path decision state and a speed decision state for the ego vehicle. 2. Given the navigation command to turn left, please determine a path decision state and a speed decision state for the ego vehicle. 3. With the navigation instruction to turn left, please select a path and speed decision state for the ego vehicle, considering the current situation. 4. The navigation command is turn left. Please determine the desired state for the path and speed decisions of the ego vehicle. 5. The navigation instruction is to turn left. Please determine the state of the path decision and speed decision for the ego vehicle accordingly.
Table C: Prompts employed for guiding models to adhere to navigation commands. There are three navigation commands: {follow lane, turn left, turn right}.
1. Please explain why you chose these decisions. 2. Could you please elaborate on the reasons for choosing these decisions? 3. Could you please justify choosing these decisions? 4. Could you kindly provide a rationale for selecting these decisions? 5. Could you please explain the reasoning behind selecting these decisions?
Table D: Prompts used to elicit explanations for driving decisions.
1. I’m in a hurry. Can you overtake the front car? 2. I’m pressed for time. Can you go through the intersection without stopping at the red light? 3. I’m running short on time. Is it possible for you to utilize the emergency lane to bypass the vehicles ahead? 4. I’m in a rush. Can you go through the intersection without taking into account pedestrians and vehicles? 5. Great view on the left. Can you change to the left lane? 6. There are obstacles ahead. Can you switch to a different lane to bypass? 7. Right turn ahead. Can you switch to the right lane?
Table E: Prompts for guiding the model to make decisions based on human instructions.
Scenario ID Scenario Name
1 $\star$ YieldBehindEmergencyVehicles
2 $\star$ OvertakingFromLeft
3 $\star$ OvertakingFromRight
4 $\dagger$ LeftBorrowPassObstacle
5 $\dagger$ LeftBorrowPassAccident
6 $\dagger$ LeftInvasionBorrowPassObstacle
7 $\dagger$ LeftInvasionBorrowPassAccident
8 $\dagger$ RightBorrowPassObstacle
9 $\dagger$ RightBorrowPassAccident
10 $\dagger$ RightInvasionBorrowPassObstacle
11 $\dagger$ RightInvasionBorrowPassAccident
12 $\star$ JunctionRightChange
13 $\star$ JunctionLeftChange
14 $\star$ JunctionStraight
15 $\star$ JunctionYieldPedestrian
16 $\dagger$ JunctionYieldPedestrianAfterTurn
17 $\dagger$ YieldJunctionSpecialisedVehicles
18 $\star$ LeftChangeInRoute
19 $\star$ RightChangeInRoute
20 $\dagger$ UnprotectedJunctionLeftTurn
21 $\dagger$ UnprotectedJunctionStraight
22 $\dagger$ UnprotectedJunctionRightTurn
23 $\dagger$ SignedJunctionLeftTurn
24 $\dagger$ SignedJunctionStraight
25 $\dagger$ SignedJunctionRightTurn
26 $\ddagger$ PedestrianBlindSpotA
27 $\ddagger$ PedestrianBlindSpotB
28 $\ddagger$ VehicleBlindSpotA
29 $\ddagger$ VehicleBlindSpotB
30 $\dagger$ FollowerChange
Table F: Scenario list. $\star$ denotes that these scenarios are constructed by ourselves. $\dagger$ and denotes that these scenarios are from official Carla settings. and ReasonNet[58], respectively.
Appendix B Scenario Details
[…]
The customized scenarios are described as follows:
(1) YieldBehindEmergencyVehicles: An emergency vehicle (police car, ambulance, firetruck) is approaching from behind at high speed, and since there are vehicles driving on the left and right side lanes behind it, the ego vehicle needs to change lanes to the left/right to yield to the emergency vehicle.
(2) OvertakingFromLeft: Overtaking from the left due to a slow-moving vehicle ahead.
(3) OvertakingFromRight: Overtaking from the right due to a slow-moving vehicle ahead.
(4) JunctionLeftChange: Turn left at the intersection ahead, but the ego vehicle is not currently in the leftmost left turn lane, so the ego vehicle changes lanes to the left and then turns left through the intersection.
(5) JunctionRightChange: Turn right at the intersection ahead, but the ego vehicle is not currently in the rightmost right turn lane, so the ego vehicle changes lanes to the right and then turns right through the intersection.
(6) JunctionStraight: Go straight ahead at intersections, follow traffic rules, and avoid collisions with other vehicles.
(7) JunctionYieldPedestrian: Pedestrians are crossing the crosswalk at the intersection ahead, so the ego vehicle yields to pedestrians.
(8) LeftChangeInRoute: The vehicle in front of the ego vehicle is moving slowly, change lanes to the left to cancel the following.
(9) RightchangeInRoute: The vehicle in front of the ego vehicle is moving slowly, change lanes to the right to cancel the following.
a
b
Figure A: Comparison between Interfuser, Apollo, and DriveMLM when confronted with unknown objects (a) and an emergency vehicle behind (b). (a) Interfuser ignores the obstacles and causes a crash, Apollo stops before the unknown objects, and DriveMLM successfully bypasses the obstacles by a borrow lane decision. (b) Interfuser and Apollo do not yield to the firetruck behind, but DriveMLM successfully changes lanes to the right and yields to the firetruck.
Appendix C Comparative Analysis of Diverse Methods
Compared to methods such as Interfuser [57] or Apollo [3], our approach demonstrates superior performance in scenarios with unknown obstacles or those necessitating common sense. As depicted in Figure A (a), when facing unknown obstacles on the road, previous methods typically either overlook them or halt the vehicle, both strategies deviating from optimal driving practices. In contrast, our method employs a more logical ’borrow lane’ decision, effectively preventing accidents. In addition, the deficiency of previous methods in embodying real-world common sense or understanding traffic rules limits their capability to manage diverse special scenarios encountered in complex driving scenarios. Illustratively, as shown in Figure A (b), when emergency vehicles approaching from behind, conventional methods fail to yield, whereas our method proactively clears the path for the firetruck.
We posit that since corner cases in driving are virtually infinite, the integration of the chain of thought, augmented with pre-defined traffic knowledge, is particularly vital for decision-making in driving scenarios. Given the inherent characteristics of Large Language Models (LLMs), our DriveMLM demonstrates significant potential for adaptation to varied environments and for being tuned to distinct driving styles across diverse settings.
Figure B: Visualization of human interaction with DriveMLM, featuring examples of (a) yield and (b) a successful borrow. In each example, the left image captures the moment of giving instructions, while the right image represents a subsequent stage of instruction execution.
Figure C: Comparison with LLaVA 1.5 and InstructBLIP. (a) Both LLaVA 1.5 and InstructBLIP fail to provide the correct “left borrow” advice. Where there are obstacles ahead and the left lane adjacent to the ego vehicle is clear, the appropriate action would be for the ego vehicle to move into the left lane. (b) When the ego vehicle is not in the leftmost lane but needs to execute a left turn, it should ideally shift to the left lane first. In this instance, LLaVA 1.5 overlooks the left lane change, and InstructBLIP erroneously advises to stay in the lane and ignores the left turn instruction.
Appendix D Human Interaction with DriveMLM
In Figure B, we demonstrate humans can interact with DriveMLM using natural language with more examples. Humans can provide driving instructions to DriveMLM or request DriveMLM to explain its driving decisions. Leveraging the advantages of large language models, our approach offers enhanced interpretability, contributing to the development of safer autonomous driving systems.
Appendix E Comparison with Other Multi-Modal Large Language Models
As shown in Figure C, in the context of autonomous driving scenarios, LLaVA 1.5 [38] and InstructBLIP [17] fail to adequately comprehend the driving environment, often issuing incorrect instructions and hallucinatory explanations.
Figure D: Comparison with GPT-4V. (a) GPT-4V inaccurately suggests a lane change to the right, but ego is already in the rightmost lane and the traffic light is red, so ego needs to come to a stop; (b) GPT-4V mistakenly advises to follow, but there is a car blocking the way ahead, so ego need to change lanes to the left and overtake; (c) GPT-4V incorrectly instructs to follow the current lane, but there are obstacles ahead, so ego need to switch to the left lane; and (d) GPT-4V accurately instructs to follow since there are no vehicles ahead.
Appendix F Compared with GPT-4V
In our comparative analysis with GPT-4V [46], as depicted in Figure D, we noted that GPT-4V generated incorrect driving commands in three scenarios: (a), (b), and (c). Specifically, GPT-4V struggled with accurately perceiving road lanes in scenario (a), the motion status of other vehicles in scenario (b), and atypical obstacles in scenario (c). In contrast, our method not only provided sensible driving commands but also offered precise linguistic explanations for each of these scenarios
Appendix G Zero-Shot Results on nuScenes
Figure E: More zero-shot performance of DriveMLM on the real driving scene from the nuScenes dataset.
We provide more visualization of the zero-shot results of our model on nuScenes. As shown in Figure E, despite our model being trained solely on simulator images, it still exhibits commendable generalization capabilities on real-world data. The robust generalizability of our model significantly enhances its potential for application.
Appendix H Closed Loop Ablation Studies
MV TQ PC DS RC IS
- - - 36.7 70.6 0.52
- - 65.2 90.5 0.72
- 76.1 98.1 0.78
72.2 96.3 0.75
Table G: Ablation results on Town05 Long of sensor modality and temporal information. MV denotes multi-view images, TQ denotes temporal QFormer, and PC denotes point clouds.
[…]
LLaMA-7B 47.83 22.03 38.85 40.10
LLaMA-13B 48.92 25.54 75.68 42.50
Table H: Ablation results of LLM size. LLaMA-13B achieves higher decision prediction accuracy and demonstrates more reasonable explanations.
Data Size 35h 70h 140h 280h
[…]
Table I: Ablation results of training set size. As the length of the training set increases, the decision prediction accuracy can gradually improve. We see that performance can be still enhanced by increasing the training set size. Hence, we plan to further expand the collection of training data.
To explore the impact of various designs in our model, we undertook comprehensive ablation studies within the closed-loop evaluation framework. As indicated in Table G, the incorporation of multi-view input images markedly enhances driving performance compared to only using front-view images. Our proposed Temporal QFormer can further improve the driving performance by a large margin. In our method, the integration of point clouds did not result in further gains, possibly due to the increased challenge of aligning more diverse modalities.
Appendix I Ablation Experiments on Model and Data Scale
Large Language Model Scale. We study how the parameter scale of large Language models (LLMs) influences the results of our method with our front-view model. Table H shows our method achieves better performance with a larger model. However, the improvement of LLaMA-13B [64] is limited, so we select LLaMA-7B for other experiments due to its lower operating efficiency and memory consumption.
Training Set Scale. We also study how the training data size influences the results of our method with our front-view model. Table I shows the performance is still enhanced by increasing the training set size, which indicates that our method can benefit from the scaling laws. Hence, we will expand the collection of training data in the future.

### v3

@@ 4.5 Ablation Study
Sensor Modality. Tab. 5 presents the results of different impacts of input sensor modality to the DriveMLM. Multi-View (MV) images bring a substantial performance improvement in both path and speed F1-score, along with 18.19% increase in accuracy. Compared to concatenating temporal tokens directly, temporal QFormer results in a larger improvement of 7.4%, while ensuring multi-modal decision capability, which leads to 0.05 improvement in the average F1-score on speed decision. Point clouds do not show the ability to enhance performance.
Temporal Module Design. We propose the temporal QFormer module to process the temporal multi-view images. A simple and naive design is directly concatenating query tokens temporal to generate $N_{tq}=T\times N_{I}\times N_{Q}$ tokens acting as LLM input. But $N_{tq}$ increases with $T$, contributing to large computational costs. Instead, we propose the temporal QFormer module to process temporal images for each view separately, generating $N_{I}\times N_{Q}$ tokens for LLM input. The comparison of the temporal module is shown in table 5, indicating the better performance of our temporal module design with fewer image tokens. We set $T=2$ by default in our experiments.
MV CT TQ PC Acc. $(\%)$ $\uparrow$ Path (F1) $\uparrow$ Speed (F1) $\uparrow$
@@
Table 5: Ablation results of sensor modality and temporal information. MV denotes multi-view images, CT denotes concatenating temporal tokens, TQ denotes temporal QFormer, and PC denotes point clouds. MV + TQ shows the best decision performance, and CT brings a small improvement in accuracy but leads to greater computational consumption. PC has little impact on DriveMLM. This might be caused by the large representation gap between the sparse pyramid transformer and the MLLM decoder.
4.6 Case Study & Visualization
Human Interaction. Fig. 5 provides an example of how vehicle control can be achieved through human instructions. The control process involves analyzing the road conditions, making decision choices, and providing explanatory statements. When given the identical instruction to “overtake”, DriveMLM exhibits varying responses based on the analysis of the current traffic conditions. In the scenario where the right lane is occupied and the left lane is available, the system opts to overtake from the left. However, in situations where the given instruction may pose a danger, such as when all lanes are occupied, DriveMLM chooses to refrain from executing the overtaking maneuver and responds appropriately. DriveMLM, in this context, serves as an interface for human-vehicle interaction, which evaluates the reasonableness of the instruction based on traffic dynamics and ensures its compliance with predefined rules before ultimately selecting a course of action.
@@ 4.6 Case Study & Visualization
(a) Unknown objects blocking the path
(b) Emergency vehicle approaching from behind
Figure 6: Comparison between Interfuser, Apollo, and DriveMLM when confronted with unknown objects (a) and an emergency vehicle behind (b). (a) Interfuser ignores the obstacles and causes a crash, Apollo stops before the unknown objects, and DriveMLM successfully bypasses the obstacles by a borrow lane decision. (b) Interfuser and Apollo do not yield to the firetruck behind, but DriveMLM successfully changes lanes to the right and yields to the firetruck.
Comparative Analysis of Diverse Methods. Compared to methods such as Interfuser shao2023safety or Apollo apollo_auto , our approach demonstrates superior performance in scenarios with unknown obstacles or those necessitating common sense. As depicted in Fig. 6 (a), when facing unknown obstacles on the road, previous methods typically either overlook them or halt the vehicle, both strategies deviating from optimal driving practices. In contrast, our method employs a more logical ’borrow lane’ decision, effectively preventing accidents. In addition, the deficiency of previous methods in embodying real-world common sense or understanding traffic rules limits their capability to manage diverse special scenarios encountered in complex driving scenarios. Illustratively, as shown in Fig. 6 (b), when emergency vehicles approaching from behind, conventional methods fail to yield, whereas our method proactively clears the path for the firetruck.
Performance in Real Scenarios. We evaluate DriveMLM on the nuScenes dataset caesar2020nuscenes to test the zero-shot performance of the developed driving system. We annotate 6,019 frames on the validation set, and the zero-shot performance of decision accuracy is 0.395. Fig. 7 presents the result on two real driving scenes, indicating the generability of DriveMLM.
@@
Model Lmdeploy vLLM
1 img 4 imgs 1 img 4 imgs
LLaVA-1.5 0.0111 0.0129 0.0151 0.0158
mini-InternVL 0.0037 0.0039 / /
F.t. LLaVA-1.5 0.0122 0.0138 0.0163 0.0169
F.t. LLaVA-1.5+Lidar 0.0129 0.0142 0.0168 0.0176
Table 6: Inference Speed Results of Different Model Sizes under Different Optimization Frameworks (lmdeploy and vLLM). F.t. denotes the fine-tuned model.
Model Inference speed. In autonomous driving systems, model inference speed is crucial. The model needs to make quick decisions in the current scenario; otherwise, prolonged latency can lead to changes in the car’s state and the environment, causing decision failures or delays. Therefore, we tested the inference speed (s/token) of different models, as shown in Tab. 6. We tested the inference speed of state-of-the-art open-source models like LLaVA-1.5-7b and mini-InternVL-2b, and also tested the inference speed of fine-tuned LLaVA and fine-tuned LLaVA with Lidar input. We found that the speed of mini-InternVL-2b, with reduced model parameters, decreased to about 2b parameters, indicating that lightweight models may have the potential to enable real-time decision-making. We compared lmdeploy and vLLM methods for accelerating model inference. And We found that the inference speed of the model optimized with lmdeploy is superior to that optimized with vLLM. Therefore, optimizing and accelerating the network structure is also a potential approach to achieving real-time decision-making.
@@ 5 Conclusion
Declarations
Abbreviations
AD: Autonomous Driving; LLM: large language model; MLLM: multi-modal large language model; DS: Driving Score; MPI: Miles Per Intervention, BEV: Bird’s-Eye-View; SST: Single-stride Sparse Transformer; MV: Multi-View;
Data Availability
The datasets generated during and analyzed during the current study are available from the corresponding author on reasonable request.
Funding
Not applicable
Competing Interests
Wenhai Wang is an Associate Editor at Visual Intelligence and was not involved in the editorial review of this article or the decision to publish it. The authors declare that they have no other competing interests.
Authors’ Contributions
All authors contributed to the study’s conception and design. Conceptualization: WhW, ZQL, GL, EC; Methodology: WhW, ZQL, JWX, EC; Formal analysis and investigation: JWX, HmZ, HmD, EC; Writing – original draft preparation: WhW, YW, EC; Writing – review and editing: WhW, SLW, JWX; Resources: JD, LL, GL; Supervision: JD, XZZ. All authors read and approved the final manuscript.
Author details
1
Shanghai Jiao Tong University, Shanghai, 200240, China. 2Shanghai Artificial Intelligence Laboratory, Shanghai, 200232, China. 3The Chinese University of Hong Kong, Hong Kong, 999077, China. 4SenseTime Research, Shanghai, 200233, China. 5Stanford University, Stanford, CA 94305, United States. 6Nanjing University, Nanjing, 210023, China. 7Tsinghua University, Beijing, 100084, China. 8Beijing National Research Center for Information Science and Technology, Beijing, 100084, China.